OpenStudy (anonymous):

Write the point-slope equation of the line through (-2, 1) with a slope of 3.

OpenStudy (hoblos):

the general form is y-y1=m(x-x1 where m is the slope and (x1,y1) is a point on the line y-1 = 3(x+2)